OpenStudy (anonymous):

Prove the Identity. sinx[sinx+sin(5x)]=cos(2x)[cos(2x)-cos(4x)]

OpenStudy (anonymous):

The Identities I am using are: \[\large sinA + sinB = 2 \cdot \sin(\frac{A + B}{2})\cos(\frac{A - B}{2})\] \[\large 2sinAsinB = \cos(A-B) - \cos(A+B)\]

OpenStudy (anonymous):

So solve the brackets first using first Identity I have given above: \[\sin(x)(\sin(x) + \sin(5x)) \implies \sin(x)( \color{green}{2 \cdot \sin(3x) \cdot \cos(2x))}\] \[\implies \color{blue} {(2 \cdot \sin(3x) \cdot \sin(x))} \cdot \cos(2x) \implies \color{cyan}{(\cos(2x) - \cos(4x))} \cdot \cos(2x)\] Rearrange them to get equal to RHS: \[\implies \color{red}{\cos(2x) } \color {darkgreen} {(\cos(2x) - \cos(4x))}\]

OpenStudy (anonymous):

@rosiepose99 go through the steps and tell me do you have any problem in any step??

OpenStudy (anonymous):

ok can you explain how you did the final step where you had (2*sin(3x)*sin(x))*cos (2x)

OpenStudy (anonymous):

why did you put the extra sin(x) before the cos(2x)

OpenStudy (anonymous):

It is not the extra sin(x) see you have sin(x) in outside too.. So I have just taken that backwards..

OpenStudy (anonymous):

in mutiplication you can change the preference order: see: \[\large 5 \times 4 \times 2 \implies 4 \times 5 \times 2 \implies 2 \times 4 \times 5\] They all are one and same thing..

OpenStudy (anonymous):

So: \[\sin(x) \times 2 \times \sin(3x) \times \cos(2x)\] I have written this as: \[2 \times \sin(3x) \times \sin(x) \times \cos(2x)\] Getting??

OpenStudy (anonymous):

I am expecting a reply from you @rosiepose99

OpenStudy (anonymous):

sorry and thank you, I understand it.